Here is another one I need help with. Thanks in Advance.
1. If I am surveying the constituents for my favorite candidate ... and I wish to be 95% confident of the estimate I develop for the proportion of the constituents that favor my candidate and i want the error to be within 2% (0.02) of the actual population proportion

>> how many folks do there need to be in my sample ??

Answer by stanbon(75887)

If I am surveying the constituents for my favorite candidate ... and I wish to be 95% confident of the estimate I develop for the proportion of the constituents that favor my candidate and i want the error to be within 2% (0.02) of the actual population proportion
>> how many folks do there need to be in my sample ??
------------------------------
Formula: E = z*sqrt(pq/n)
sqrt(n) = z*(sqrt(pq)}/E
n = [z/E}^2*pq
------------------
z-score for 95^ confidence = 1.96
Since p is not given, avoid bias by letting p=1/2=q
E is given as 0.02
-------------------
n = [1.96/0.02]^2(1/4)
n = 9604*(1/4)
n = 2401 (desired sample size)
===============================
Cheers,
Stan H.